OpenStudy (anonymous):

4x/x^2 - 9 all over 8x^2/x^2 + 6x + 9

OpenStudy (anonymous):

I don't see the simplifies result @MarkRijckenberg

OpenStudy (anonymous):

Is this the expression you're looking at? \[\frac{\frac{4x}{x^2 - 9}}{\frac{8x^2}{x^2 + 6x + 9}}\]

OpenStudy (anonymous):

It's a bit ambiguous when written out horizontally like that. I recommend liberal use of parentheses.

OpenStudy (anonymous):

Yes that is correct.

OpenStudy (anonymous):

If you're going to enter it into a calculator (or wolframalpha), then you'll need to express it like this: (4x/(x^2 - 9)) / ((8x^2)/(x^2 + 6x + 9)) Even then, I tried that in wolframalpha, and it didn't give the fully simplified result.

OpenStudy (anonymous):

\[1/2\,{\frac {x+3}{ \left( x-3 \right) x}}\]

OpenStudy (anonymous):

Best first step is to express the fraction-division as fraction-mulitplication. (Division is defined as multiplication by the reciprcal.) Then factor the crap out of everything and see what cancels.

OpenStudy (anonymous):

from wolfram if you type it like this 4*x/((x^2-9)*(8*x^2/(x^2+6*x+9)))

OpenStudy (anonymous):

Or you could just write it down and do the algebra . . .

OpenStudy (markrijckenberg):

Guys, both your formula's give the same results in WolfRam....

OpenStudy (anonymous):

Yep, and neither shows Lindsey how to do it herself.

OpenStudy (anonymous):

Intermediate step looks like this: \[\frac{4x}{(x+3)(x-3)} \space \times \space \frac{(x+3)(x+3)}{8x^2}\]